# Theory of Photocurrent and High-Harmonic Generation with Chiral Fermions

### v1

@@ Abstract
$I_{\parallel}(n)=0\quad\mathrm{except~{}for}\quad n\equiv\pm 1\pmod{N},$ (9)
$I_{\perp}(n)=0\quad\mathrm{except~{}for}\quad n\equiv 0\pmod{N}.$ (10)
@@
The nonlinear optical responses of B20-type chiral semimetals have attracted significant interest, with several experimental groups investigating the detection of the circular photogalvanic effect Rees et al. (2020); Ni et al. (2020, 2021). Recent advances in high-intensity laser technology have facilitated experimental studies on the nonlinear optical responses of topological materials Wu et al. (2017); Patankar et al. (2018); Osterhoudt et al. (2019); Sirica et al. (2019). In particular, HHG in topological insulators and Weyl semimetals has been experimentally observed by several groups at electric field strengths comparable to or exceeding those considered in this work Bai et al. (2021); Schmid et al. (2021); Lv et al. (2021); Baykusheva et al. (2021); Heide et al. (2022). As proposed in this study, measuring the dc photocurrent beyond the second-order regime and observing parity-dependent directivity of HHG provide valuable opportunities to explore the nonequilibrium properties of chiral semimetals.
It is known that incorporating spin-orbit coupling modifies the energy bands and alters the topological properties of chiral fermions Tang et al. (2017); Chang et al. (2017); Pshenay-Severin et al. (2018). Such band modifications are not expected to affect the qualitative features of the main results of this work, whereas the inclusion of spin-orbit coupling is expected to improve the accuracy of the theoretical estimation of the photocurrent, as confirmed in previous numerical simulations on optical conductivities Xu et al. (2020); Ni et al. (2020, 2021).
Conclusion.—We demonstrate the bulk dc photocurrent and high-harmonic generation (HHG) induced by circularly polarized light in B20-type chiral semimetals. The dc photocurrent exhibits multiple sign changes depending on the light electric field $E$ and the frequency $\Omega$, reflecting the nonreciprocal excitation of chiral fermions and the underlying chirality of the lattice structure. For HHG, the directivity of the HHG intensity depends on the parity of the harmonic orders. Namely, the HHG intensity of odd-order harmonics exhibits directivity perpendicular relative to the polarization plane, while that of even-order harmonics exhibits directivity within the polarization plane. The parity-dependent directivity of HHG, along with the appearance of even-harmonic generation, is the distinctive feature of chiral semimetals. This study provides a foundation for exploring nonlinear optical responses induced by circularly polarized light and offers insights into the nonequilibrium properties of chiral semimetals, paving the way for future experimental and theoretical advancements in the strongly driven chiral systems.
Acknowledgments.— The authors are grateful to A. Yamakage for valuable discussions. This work was supported by JSPS KAKENHI (Grants No. JP20H00337, No. JP23H04522 and No. JP24H02231), JST CREST (Grant No. JPMJCR20T1), and Waseda University Grant for Special Research Projects (Grant No. 2024C-153).

### v2

@@ IV Discussion
$I_{\parallel}(n)=0\quad\mathrm{except\leavevmode\nobreak\ for}\quad n\equiv\pm 1\pmod{N},$ (9)
$I_{\perp}(n)=0\quad\mathrm{except\leavevmode\nobreak\ for}\quad n\equiv 0\pmod{N}.$ (10)
@@
The nonlinear optical responses of B20-type chiral semimetals have attracted significant interest, with several experimental groups investigating the detection of the circular photogalvanic effect [58, 59, 60]. Recent advances in high-intensity laser technology have facilitated experimental studies on the nonlinear optical responses of topological materials [79, 25, 14, 15]. In particular, HHG in topological insulators and Weyl semimetals has been experimentally observed by several groups at electric field strengths comparable to or exceeding those considered in this work [29, 30, 26, 31, 80]. As proposed in this study, measuring the dc photocurrent beyond the second-order regime and observing parity-dependent directivity of HHG provide valuable opportunities to explore the nonequilibrium properties of chiral semimetals.
It is known that incorporating spin-orbit coupling modifies the energy bands and alters the topological properties of chiral fermions [47, 49, 48]. Such band modifications are not expected to affect the qualitative features of the main results of this work, whereas the inclusion of spin-orbit coupling is expected to improve the accuracy of the theoretical estimation of the photocurrent, as confirmed in previous numerical simulations on optical conductivities [50, 59, 60].
V Conclusion
We demonstrate the bulk dc photocurrent and high-harmonic generation (HHG) induced by circularly polarized light in B20-type chiral semimetals. The dc photocurrent exhibits multiple sign changes depending on the light electric field $E$ and the frequency $\Omega$, reflecting the nonreciprocal excitation of chiral fermions and the underlying chirality of the lattice structure. For HHG, the directivity of the HHG intensity depends on the parity of the harmonic orders. Namely, the HHG intensity of odd-order harmonics exhibits directivity perpendicular relative to the polarization plane, while that of even-order harmonics exhibits directivity within the polarization plane. The parity-dependent directivity of HHG, along with the appearance of even-harmonic generation, is the distinctive feature of chiral semimetals. This study provides a foundation for exploring nonlinear optical responses induced by circularly polarized light and offers insights into the nonequilibrium properties of chiral semimetals, paving the way for future experimental and theoretical advancements in the strongly driven chiral systems.
Acknowledgments
Y.O. is grateful to A. Yamakage for valuable discussions. This work was supported by JSPS KAKENHI (Grants No. JP20H00337, No. JP24H02231, and No. JP25H00611), JST CREST (Grant No. JPMJCR20T1), and Waseda University Grant for Special Research Projects (Grants No. 2024C-153, No. 2024C-290, and No. 2025C-133).